OpenStudy (anonymous):

The scale of the map is 1:25000. The actual area of the school is 25000m^2. Calculate, in cm^2, the area represented on the map

OpenStudy (anonymous):

25000 m^2 = 250000000 cm^2 area= 250000000/25000 = 10000 cm^2